SPICA Online User's manual

Weather conditions and cameras at Mount Wilson

The CHARA Array wiki

Quality Check SPICA (in progress)


UT 2-6 May (on site, Denis, Juraj, Mathieu, Pierre)
  • pdfMay 1: Interesting engineering tests on E1-AO and different synthetic matrices.
  • May 2: Fog, rain and humidity. Lot of daytime activity on AO, alignment, dispersion, injection, spectral calibration
  • May 3: Fog, humidity - no obs. Daytime activities on AO, alignment procedure, and issue probably found for the correction of dispersion.
  • May 4: 100% humidity, Winter advisory, snow expected. No observations, continuation of daytime activities
  • May 5: A few inches of snow during the day, fog and humidity. Continuation of daytime activities.
  • May 6
UT 6-10 April:(engineering, remote Denis&Fred)
  • On April 6 & 7, CHARA is still closed after the winter break
  • pdfApril 8: clouds for the first half of the time, then S1S2 fringes
  • pdfApril 9: humidity for the first halp of the night then more fringes
  • pdfApril 10: humidity for the first half of the night then alignment, AO, and more fringes
UT 1-5 March: On site operation (Denis, Pierre, Daniel) but observatory closed because of snow storm


UT 1-5 December: Remote operations but no on-sky fringes because of poor weather


UT 4-8 November: Remote operations (Cyril, Sylvain, Fred, Nicolas, Denis + Julien)

 Engineering activities on LDC/VLDC/DDL - Offsets on sky - Fringe tracking

pdfLog20221104.pdf pdfLog20221105.pdf pdfLog20221106.pdf

W2W1 fringes seen on the spicaRTD after stretching but still with some dispersion: movIMG_2724.MOV


UT 4-11 October: onsite operations (Pierre, Julien, Denis)

More engineering activities on LDC/VLDC/DDL - Offsets on sky - AO & Fringe tracking - First set of initial data

pdf20221004.pdf pdf20221005.pdf pdf20221006.pdf  pdf20221007.pdf  pdf20221008.pdf  pdf20221009.pdf  pdf20221010.pdf  pdf20221011.pdf

 S1S2 fringes on gam Cas as seen on the RTD (dispersion not completely compensated) movIMG_2691.MOV


UT 12-19 August: Remote from Calern

Engineering activities on dispersion, baseline solution, and operations.

UT 8-12 July 2022: Remote from Calern

UT 21-25 June 2022: Mt Wilson (Denis, Fred, Julien)

  • 25 June 2022: Commissioning, MLDC only
    • Fringes S1S2 obtained on different stars, always the same offset (==> beam samplers different from STS)
    • Fringes W1W2 on HD177724 at +3850µm on W1. Again a BS/STS issue.
    • New TELAO telemetries under very good r0 conditions (18cm)
  • 24 June 2022: Commissioning, but poor seeing
    • New records of control and science data for tests of injection
    • Hard time to find fringes because of a bad settings of the MLDC+VLDC
    • New record of TELAO telemetries
  • 23 June 2022: Array closed following the power outage
  • 22 June 2022: High winds, thunderstorms
  • 21 June 2022: Commissioning (first 2T fringes)
    • S1S2 fringes found on Eta Uma, MR720 and LR - Important dispersion
    • Recording of Control and Science data
    • Recording of sequence of injection (control and science)
    • Recording of HD148857 (binary with separation 1.3'') for field characterization
    • TELAO telemetries recorded
    • Control data recorded with different settings of LABAO focus on S1/B5

UT 27-31 May 2022: Nice&Calern (Sylvain, Cyril, Philippe, Fred, Denis)

  • 31 May 2022: spica-vis software activities
  • 30 May 2022: spica-vis software activities
  • 29 May 2022: very poor seing conditions, no SPICA-FT close loop possible
  • 28 May 2022: operation of SPICA-FT, GD and PD loops up to 4 telescopes
  • 27 May 2022: clouds but promising tests of SPICA-FT control loops

UT 23-27 April 2022: Mt Wilson (David, Denis, Fred)

  • 27 April 2022: clouds and poor seeing.
  • 26 April 2022: medium conditions. Tests of injection
  • 25 April 2022: good conditions. Validation of the FTT and fiberexplorer but still low ejection level. Tests of refraction.
  • 24 April 2022: very bad seeing, tests of alignment procedures
  • 23 April 2022: clouds and hgh humidity

UT 28 Mars to 31 March 2022: Calern (Denis, Fred, Philippe, Sylvain)

  • 31 March 2022: poor seeing, high humidity. Continuation of software activities and SPICA-FT closed loop.
  • 30 MArch 2022: high humidity, no observations
  • 29 March 2022: bad weather, no observations
  • 28 March 2022: bad weather, no observations

UT 10-12 March 2022: Mt Wilson (Cyril, Daniel, Denis, Julien, Philippe, Pierre)

  • 12 March 2022: tests of AO tuning and fiber injection.
  • 11 March 2022: poor seeing and winds, no observations
  • 10 March 2022: (First stellar light) alignment on sky; very poor seeing. Issues for injection and fringes.